OpenStudy (anonymous):

How you you graph y = -2x + 1 ?

jimthompson5910 (jim_thompson5910):

plug in x = 0 and evaluate to find y what do you get?

jimthompson5910 (jim_thompson5910):

y = -2x + 1 y = -2*0 + 1 ... plug in x = 0 y = ???

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

so you have one point (0, 1)

jimthompson5910 (jim_thompson5910):

plug in x = 1 y = -2x + 1 y = -2*1 + 2 y = ???

OpenStudy (anonymous):

ohhhhhhhhhhhh , I get the drift.

jimthompson5910 (jim_thompson5910):

so what do you get for y ?

OpenStudy (anonymous):

0 right ?

jimthompson5910 (jim_thompson5910):

so you have the two points (0, 1) and (1, 0)

jimthompson5910 (jim_thompson5910):

oh i made a typo, i just realized that...sry

jimthompson5910 (jim_thompson5910):

plug in x = 1 y = -2x + 1 y = -2*1 + 1 y = -2+1 y = -1 So when x = 1, y is y = -1 Therefore, we have the point (1, -1)

jimthompson5910 (jim_thompson5910):

so what you really have are these two points (0,1) and (1, -1) Plot them both. Then draw a straight line through them to graph y = -2x+1

OpenStudy (anonymous):

what about y = -2/3 x + 1